### v1

# Disentanglement in a GAN for Unconditional Speech Synthesis

Abstract
Can we develop a model that can synthesize realistic speech directly from a latent space, without explicit conditioning? Despite several efforts over the last decade, previous adversarial and diffusion-based approaches still struggle to achieve this, even on small-vocabulary datasets. To address this, we propose AudioStyleGAN (ASGAN) – a generative adversarial network for unconditional speech synthesis tailored to learn a disentangled latent space. Building upon the StyleGAN family of image synthesis models, ASGAN maps sampled noise to a disentangled latent vector which is then mapped to a sequence of audio features so that signal aliasing is suppressed at every layer. To successfully train ASGAN, we introduce a number of new techniques, including a modification to adaptive discriminator augmentation which probabilistically skips discriminator updates. We apply it on the small-vocabulary Google Speech Commands digits dataset, where it achieves state-of-the-art results in unconditional speech synthesis. It is also substantially faster than existing top-performing diffusion models. We confirm that ASGAN’s latent space is disentangled: we demonstrate how simple linear operations in the space can be used to perform several tasks unseen during training. Specifically, we perform evaluations in voice conversion, speech enhancement, speaker verification, and keyword classification. Our work indicates that GANs are still highly competitive in the unconditional speech synthesis landscape, and that disentangled latent spaces can be used to aid generalization to unseen tasks. Code, models, samples: https://github.com/RF5/simple-asgan/.
Index Terms:
Unconditional speech synthesis, generative adversarial networks, speech disentanglement.
publicationid: pubid: © 2023 IEEE
I Introduction
Unconditional speech synthesis systems aim to produce coherent speech without conditioning inputs such as text or speaker labels [1]. In this work we are specifically interested in learning to map noise from a known continuous distribution into spoken utterances [2]. A model that could do this would have several useful downstream applications: from latent interpolations between utterances and fine-grained tuning of properties of the generated speech, to audio compression and better probability density estimation of speech. Some of these advances from latent generative modelling have already been realized in the image modality [3, 4]; our goal is to bring these developments to the speech domain.
Direct speech synthesis from a latent space is a very challenging problem, and we therefore restrict ourselves to a limited vocabulary, as was also done in previous work [1, 5, 6]. In this problem setting, recent studies on diffusion models [7] for images [8, 9, 10] has led to major improvements in unconditional speech synthesis. The current best-performing approaches are all based on diffusion modelling [6, 5], which iteratively de-noise a sampled signal into a waveform through a Markov chain [7]. Before this, many studies used generative adversarial networks (GANs) [11] that map a latent vector to a sequence of speech features with a single forward pass through the model. However, performance was limited [1, 2], leading to GANs falling out of favour for this task.
Motivated by the recent developments in the StyleGAN literature [12, 13, 3] for image synthesis, we aim to reinvigorate GANs for unconditional speech synthesis, where we are particularly interested in their ability for learning continuous, disentangled latent spaces [13]. To this end, we propose AudioStyleGAN (ASGAN): a convolutional GAN which maps a single latent vector to a sequence of audio features, and is designed to have a disentangled latent space. The model is based in large part on StyleGAN3 [3], which we adapt for audio synthesis. Concretely, we adapt the style layers to remove signal aliasing caused by the non-linearities in the network. This is accomplished with anti-aliasing filters to ensure that the Nyquist-Shannon sampling limits are met in each layer. We also propose a modification to adaptive discriminator augmentation [14] to stabilize training by randomly dropping discriminator updates based on a guiding signal.
In unconditional speech synthesis experiments, we measure the quality and diversity of the generated samples using objective metrics. We show that ASGAN sets a new state-of-the-art in unconditional speech synthesis on the Google Speech Commands digits dataset [15]. Objective metrics that measure latent space disentanglement indicate that ASGAN has smoother latent representations compared to existing diffusion models. It not only outperforms the best existing models but is also faster to train and faster in inference. Subjective mean opinion scores (MOS) indicate that ASGAN’s generated utterances sound more natural (MOS: 3.68) than the existing best model (SaShiMi [6], MOS: 3.33). We also perform ablation experiments showing intrinsically that our proposed anti-aliasing and adaptive discriminator augmentation techniques are necessary for high-quality and diverse synthesis.
This work is an extension of the conference paper [16], where (apart from the ablation experiments) many of the above intrinsic evaluations were already presented. Here, for the first time, we also present a thorough evaluation of ASGAN’s abilities – we show that ASGAN’s disentangled latent space allows us to perform several tasks unseen during training through simple linear operations in its latent space. Concretely, we demonstrate compelling zero-shot performance on voice conversion, speech enhancement, speaker verification and keyword classification on the Google Speech Commands digits dataset. While not matching the performance of state-of-the-art task-specific systems on all these tasks, our experiments show that a single model designed for disentanglement can achieve reasonable performance across a range of tasks that it hasn’t seen in training. Our work shows the continued competitive nature of GANs compared to diffusion models, and the generalization benefits of designing for disentanglement.
The paper is organized as follows. We discuss related work in Sec. II and then go on to propose ASGAN in Sec. III. The main unconditional speech synthesis experiments and their results are given in Sec. IV and V. This is followed by the experiments on the unseen tasks that ASGAN can be used for in Sec. VI, VII, and VIII.
II Related Work
Since we focus on the proposed generalization abilities provided by a continuous latent space, we first distinguish what we call unconditional speech synthesis to the related but different task of generative spoken language modelling (GSLM) [17]. In GSLM, a large autoregressive language model is typically trained on discovered discrete units (e.g. HuBERT [18] clusters or clustered spectrogram features), similar to how a language model is trained on text [19, 20]. While this also enables the generation of speech without any conditioning input, GSLM implies a model structure consisting of an encoder to discretize speech, a language model, and a decoder to convert the discrete units back into a waveform [17]. By using discrete units, GSLM approaches can produce long sequences while retaining good performance. The downside of this discrete approach is that, during generation, you are bound by the discrete units in the model. E.g. it is not possible to interpolate between two utterances in a latent space or to directly control speaker characteristics during generation. If this is desired, additional components must be explicitly built into the model [20].
In contrast, in unconditional speech synthesis we do not assume any knowledge of particular aspects of speech beforehand. Instead of using some intermediate discretization step, such models use noise to directly generate speech, often via some latent representation. The latent space should ideally be disentangled, allowing for better control of the generated speech. In contrast to GSLM, the synthesis model should learn to disentangle without being explicitly designed to control specific speech characteristics. In some sense this is a more challenging task than GSLM, which is why most unconditional speech synthesis models are still evaluated on short utterances of isolated spoken words [1] (as we also do here). In more structured conditional synthesis tasks such as text-to-audio or text-to-video, recent studies [21, 22] have demonstrated the benefits of modelling a continuous latent space from noise, and then performing synthesis from that latent space. We aim to apply this reasoning to the unconditional speech synthesis domain in an attempt to realize similar benefits.
Within unconditional speech synthesis, a substantial body of work focuses on either autoregressive models [23] – generating a current sample based on previous outputs – or diffusion models [5]. Diffusion models iteratively denoise a sampled signal into a waveform through a Markov chain with a constant number of steps [7]. At each inference step, the original noise signal is slightly denoised until, in the last step, it resembles coherent speech. Autoregressive and diffusion models are relatively slow because they require repeated forward passes through the model during inference.
Earlier studies [1, 2] attempted to use GANs [11] for unconditional speech synthesis, which has the advantage of requiring only a single pass through the model. While results showed some initial promise, performance was poor in terms of speech quality and diversity, with the more recent diffusion models performing much better [6]. However, there have been substantial improvements in GAN-based modelling for image synthesis in the intervening years [14, 12, 13]. Our goal is to improve the performance of the earlier GAN-based unconditional speech synthesis models by adapting lessons from these recent image synthesis studies.
Some of these innovations in GANs are modality-agnostic: $R_{1}$ regularization [24] and exponential moving averaging of generator weights [25] can be directly transferred from the vision domain to speech. Other techniques, such as the carefully designed anti-aliasing filters between layers in StyleGAN3 [3] require specific adaptation; in contrast to images, there is little meaningful information in speech below 300 Hz, necessitating a redesign of the anti-aliasing filters.
In a very related research direction, Beguš et al. [26, 2, 27] and Chen and Elsner [28] have been studying how GAN-based unconditional speech synthesis models internally perform lexical and phonological learning, and how this relates to human learning. Most of these studies, however, rely on older GAN synthesis models. We hope that by developing better performing GANs for unconditional speech synthesis, we can contribute to improving such investigations. Recently, [29] attempted to directly use StyleGAN2 for conditional and unconditional synthesis of emotional vocal bursts. This further motivates a reinvestigation of GANs, but here we look specifically at the generation of speech rather than paralinguistic sounds.
III Audio Style GAN
Figure 1: The ASGAN generator $G$ (left) and discriminator $D$ (right). FF, LPF, Conv1D indicate Fourier feature [3], low-pass filter, and 1D convolution layers, respectively. The number of output features/channels are indicated above linear and convolutional layers. Stacked blocks indicate a layer repeated sequentially.
Our model is based on the StyleGAN family of models [12] for image synthesis. We adapt and extend the approach to audio, and therefore dub our model AudioStyleGAN (ASGAN).
The model follows the setup of a standard GAN with a single generator network $G$ and discriminator network $D$ [11]. The generator $G$ accepts a vector $\mathbf{z}$ sampled from a normal distribution and processes it into a sequence of speech features $X$. In this work, we restrict the sequence of speech features $X$ to always have a fixed pre-specified duration. The discriminator $D$ accepts a sequence of speech features $X$ and yields a scalar output. The model uses a non-saturating logistic loss [11] whereby $D$ is optimized to raise its output for $X$ sampled from real data and minimize its output for $X$ produced by the generator. Meanwhile, the generator $G$ is optimized to maximize $D(X)$ for $X$ sampled from the generator, i.e. when $X=G(\mathbf{z})$. The speech features $X$ are converted to a waveform using a pretrained HiFiGAN vocoder [30]. During training, a new adaptive discriminator updating technique is added to ensure stability and convergence. Each component is described in detail below.
III-A Generator
The architecture of the generator $G$ is shown on the left of Fig. 1. It consists of a latent mapping network $W$ that converts $\mathbf{z}$ to a disentangled latent space, a special Fourier feature layer which converts a single vector from this latent space into a sequence of cosine features of fixed length, and finally a convolutional encoder which iteratively refines the cosine features into the final speech features $X$.
III-A1 Mapping network
The mapping network $W$ is a simple multi-layer perceptron consisting of several linear layers with leaky ReLU activations between them. As input, it takes in a normally distributed vector $\mathbf{z}\sim Z=\mathcal{N}(\mathbf{0},\mathbf{I})$; in all our experiments we use a 512-dimensional multivariate normal vector, $\mathbf{z}\in\mathbb{R}^{512}$. Passing this vector through the mapping network produces a latent vector $\mathbf{w}=W(\mathbf{z})$ of the same dimensionality as $\mathbf{z}$. As explained in [12], the primary purpose of $W$ is learn to map noise to a linearly disentangled $W$ space, as ultimately this will allow for more controllable and understandable synthesis. $W$ is coaxed into learning such a disentangled representation because it can only linearly modulate channels of the cosine features in each layer of the convolutional encoder (see details below). So, if $\mathbf{w}$ is to linearly shape the speech features, $W$ must learn a mapping that organizes the random normal space $\mathbf{z}$ into one which linearly disentangles common factors of speech variation.
III-A2 Convolutional encoder
The convolutional encoder begins by linearly projecting $\mathbf{w}$ as the input to a Fourier feature layer [31] as shown in Fig. 1. Concretely, we use the Gaussian Fourier feature mapping from [31] and incorporate the transformation proposed in StyleGAN3 [3]. This layer samples a frequency and phase from a Gaussian distribution for each output channel (fixed at initialization). The layer then linearly projects the input vector to a vector of phases which are added to the initial random phases. The output sequence is obtained as the plot of a cosine function with these frequencies and phases, one frequency/phase for each output channel. Intuitively, the layer provides a way to learn a mapping between a single vector ($\mathbf{w}$) into a sequence of vectors at the output of the Fourier feature layer, which provides the base from which the rest of the model can iteratively operate on the feature sequence to eventually arrive at a predicted mel-spectrogram.
The sequence produced by the Fourier feature layer is iteratively passed through Style Blocks, which is based on the encoder layers of the StyleGAN family of models. In each layer, the input sequence and style vector $\mathbf{w}$ are passed through a modulated convolution layer [13]: the final convolution kernel is computed by multiplying the layer’s learnt kernel with the style vector derived from $\mathbf{w}$, broadcasted over the length of the kernel. In this way, the latent vector $\mathbf{w}$ linearly modulates the kernel in each convolution.
To ensure the signal does not experience aliasing due to the non-linearity, the leaky ReLU layers are surrounded by layers responsible for anti-aliasing (explained below). All these layers comprise a single Style Block, which is repeated in groups of 5, 4, 3, and finally 2 blocks. The last block in each group upsamples by 4×4$\times$4 × instead of 2×2$\times$2 ×, thereby increasing the sequence length by a factor of 2 for each group. A final 1D convolution layer projects the output from the last group into the audio feature space (e.g. log mel-spectrogram or HuBERT features), as illustrated in the middle of Fig. 1.
III-A3 Anti-aliasing filters
[…]
III-B Discriminator
The discriminator $D$ has a convolutional architecture similar to [13]. It takes a sequence of speech features $X$ as input and predicts whether it is generated by $G$ or sampled from the dataset. Concretely, $D$ consists of four ConvD Blocks and a network head, as show in Fig. 1. Each ConvD Block consists of 1D convolutions with skip connections, and a downsampling layer with an anti-aliasing LPF in the last skip connection. The LPF cutoff is set to the Nyquist frequency for all layers. The number of layers and channels are chosen so that $D$ has roughly the same number of parameters as $G$. $D$’s head consists of a minibatch standard deviation [25] layer and a 1D convolution layer before passing the flattened activations through a final linear projection head to arrive at the logits. Both $D$ and $G$ are trained using the non-saturating logistic loss [11].
III-C Vocoder
The generator $G$ and discriminator $D$ operate on sequences of speech features and not on raw waveform samples. Once both $G$ and $D$ are trained, we need a way at convert these speech features back to waveforms. For this we use a pretrained HiFi-GAN vocoder [30] that vocodes either log mel-scale spectrograms [32] or HuBERT features [18]. HuBERT is a self-supervised speech representation model that learns to encode speech in a 50 Hz vector sequence using a masked token prediction task. These learnt features are linearly predictive of several high-level characteristics of speech such as phone identity, making it useful as a feature extractor when trying to learn disentangled representations.
III-D Implementation
We train two variants of our model: a log mel-spectrogram model and a HuBERT model. And, extending our initial investigation in [16], we train additional variants of the HuBERT-based model to understand key design choices.
III-D1 ASGAN variants
The log mel-spectrogram model architecture is shown in Fig. 1, where mel-spectrograms are computed with 128 mel-frequency bins at a hop and window size of 10 ms and 64 ms, respectively. Each 10 ms frame of the mel-scale spectrograms is scaled by taking the natural logarithm of the spectrogram magnitude. The HuBERT-based model is identical except that it only uses only half the sequence length (since HuBERT features are 20 ms instead of the 10 ms spectrogram frames) and has a different number of output channels in the four groups of Style Blocks: [1024, 768, 512, 512] convolution channels instead of the [1024, 512, 256, 128] used for the mel-spectrogram model. This change causes the HuBERT variant to contain 51M parameters, as opposed to the 38M parameters in the mel-spectrogram model.
III-D2 Vocoder variants
The HiFi-GAN vocoder for both HuBERT and mel-spectrogram features is based on the original author’s implementation [30]. The HuBERT feature HiFi-GAN is trained on the LibriSpeech train-lean-100 multispeaker speech dataset [33] to vocode activations extracted from layer 6 of the pretrained HuBERT Base model provided with fairseq [34]. The mel-spectrogram HiFiGAN is trained on the Google Speech Commands dataset. Both are trained using the original V1 HiFi-GAN configuration (number of updates, learning and batch size parameters) from [30].
III-D3 Optimization
Both ASGAN variants are trained with Adam [35] ($\beta_{1}=0,\beta_{2}=0.99$), clipping gradient norms at 10, and a learning rate of $3\cdot 10^{-3}$ for 520k iterations with a batch size of 32. Several critical tricks are used to stabilize GAN training: (i) equalized learning rate is used for all trainable parameters [25]; (ii) leaky ReLU activations with $\alpha=0.1$; (iii) exponential moving averaging for the generator weights (for use during evaluation) [25]; (iv) $R_{1}$ regularization [24]; and (v) a 0.01-times smaller learning rate for the mapping network $W$, since it needs to be updated slower compared to the convolution layers in the main network branch [3].
III-D4 Adaptive discriminator updates
We also introduce a new technique for updating the discriminator. Concretely, we first scale $D$’s learning rate by 0.1 compared to the generator as otherwise we find it overwhelms $G$ early on in training. Additionally we employ a dynamic method for updating $D$, inspired by adaptive discriminator augmentation [14]: during each iteration, we skip $D$’s update with probability $p$. The probability $p$ is initialized at 0.1 and is updated every 16th generator step or whenever the discriminator is updated. We keep a running average $r_{t}$ of the proportion of $D$’s outputs on real data $D(X)$ that are positive (i.e. that $D$ can confidently identify as real). Then, if $r_{t}$ is greater than 0.6 we increase $p$ by 0.05 (capped at 1.0), and if $r_{t}$ is less than 0.6 we decrease $p$ by 0.05 (limited at 0.0). In this way, we adaptively skip discriminator updates. When $D$ becomes too strong, both $r_{t}$ and $p$ rise, and so $D$ is updated less frequently. Conversely, when $D$ becomes too weak, it is updated more frequently. We found this new modification to be critical for ensuring that $D$ does not overwhelm $G$ during training.
We also use the traditional adaptive discriminator augmentation [14] where we apply the following transforms to the model input with the same probability $p$: (i) adding Gaussian noise with $\sigma=0.05$; (ii) random scaling by a factor of $1\pm 0.05$; and (iii) randomly replacing a subsequence of frames from the generated speech features with a subsequence of frames taken from a real speech feature sequence. This last augmentation is based on the fake-as-real GAN method [36] and is important to prevent gradient explosion later in training.
III-D5 Anti-aliasing filters
[…]
All models are trained using mixed FP16/FP32 precision on a single NVIDIA Quadro RTX 6000 using PyTorch 1.11. Trained models and code are available at https://github.com/RF5/simple-asgan/.
IV Experimental Setup: Unconditional Speech Synthesis
IV-A Data
To compare to existing unconditional speech synthesis models, we use the Google Speech Commands dataset of isolated spoken words [15]. As in other studies [6, 5, 1], we use the subset corresponding to the ten spoken digits “zero” to “nine” (called SC09). The digits are spoken by various speakers under different channel conditions. This makes it a challenging benchmark for unconditional speech synthesis. All utterances are roughly a second long and are sampled at 16 kHz, where utterances less than a second are padded to a full second.
IV-B Evaluation metrics
We train and validate our models on the official training/validation/test split from SC09. We then evaluate unconditional speech synthesis quality by seeing how well newly generated utterances match the distribution of the SC09 test split. We use metrics similar to those for image synthesis; they try to measure either the quality of generated utterances (realism compared to actual audio in the test set), or the diversity of generated utterances (how varied the utterances are relative to the test set), or a combination of both.
These metrics require extracting features or predictions from a supervised speech classifier network trained to classify the utterances from SC09 by what digit is spoken. While there is no consistent pretrained classifier used for this purpose, we opt to use a ResNeXT architecture [38], similar to previous studies [6, 5]. The trained model has a 98.1% word classification accuracy on the SC09 test set, and we make the model code and checkpoints available for future comparisons. [^1] https://github.com/RF5/simple-speech-commands Using either the classification output or 1024-dimensional features extracted from the penultimate layer in the classifier, we consider the following metrics.
  • Inception score (IS) measures the diversity and quality of generated samples by evaluating the Kullback-Leibler (KL) divergence between the label distribution from the classifier output and the mean label distribution over a set of generated utterances [39].
  • Modified Inception score (mIS) extends the diversity measurement aspect of IS by incorporating a measure of intra-class diversity (in our case over the ten digits) to reward models with higher intra-class entropy [40].
  • Fréchet Inception distance (FID) computes a measure of how well the distribution of generated utterances matches the test-set utterances by comparing the classifier features of generated and real data [41].
  • Activation maximization (AM) measures generator quality by comparing the KL divergence between the classifier class probabilities from real and generated data, while penalizing high classifier entropy samples produced by the generator [42]. Intuitively, this attempts to account for possible class imbalances in the training set and intra-class diversity by incorporating a term for the entropy of the classifier outputs for generated samples.
A major motivation for ASGAN’s design is latent-space disentanglement. In the experiments proceeding from Sec. VIII and onwards, we show this property allows the model to be applied to extrinsic tasks that it is not explicitly trained for. But before this (Sec. V), we intrinsically evaluate disentanglement using two metrics on the $Z$ and $W$ latent spaces.
  • Path length measures the mean $L_{2}$ distance moved by the classifier features when the latent point ($\mathbf{z}$ or $\mathbf{w}$) is randomly perturbed slightly, averaged over many perturbations [12]. A lower value indicates a smoother latent space.
  • Linear separability utilizes a linear support vector machine (SVM) to classify the digit of a latent point. The metric is computed as the additional information (in terms of mean entropy) necessary to correctly classify an utterance given the class prediction from the linear SVM [12]. A lower value indicates a more linearly disentangled latent space.
These metrics are averaged over 5000 generated utterances for each model. As in [12], for linear separability we exclude half the generated utterances for which the ResNeXT classifier is least confident in its prediction.
To give an indication of naturalness, we compute an estimated mean opinion score (eMOS) using a pretrained Wav2Vec2 small baseline from the VoiceMOS challenge [43]. This model is trained to predict the naturalness score that a human would assign to an utterance from 1 (least natural) to 5 (most natural). We also perform an actual subjective MOS evaluation using Amazon Mechanical Turk to obtain 240 opinion scores for each model with 12 speakers listening to each utterance. Finally, the speed of each model is also evaluated to highlight the benefit that GANs can produce utterances in a single inference call, as opposed to the many inference calls necessary with autoregressive or diffusion models.
IV-C Baseline systems
We compare to the following unconditional speech synthesis methods (Sec. II): WaveGAN [1], DiffWave [5], autoregressive SaShiMi and Sashimi+DiffWave [6]. Last-mentioned is the current best-performing model on SC09. For WaveGAN we use the trained model provided by the authors [1], while for DiffWave we use an open-source pretrained model. [^2] https://github.com/RF5/DiffWave-unconditional For the autoregressive SaShiMi model, we use the code provided by the authors to train an unconditional SaShiMi model on SC09 for 1.1M updates [6]. [^3] https://github.com/RF5/simple-sashimi Finally, for the SaShiMi+DiffWave diffusion model, we modify the autoregressive SaShiMi code and combine it with DiffWave according to [6]; we train it on SC09 for 800k updates with the hyperparameters in the original paper [6].33footnotemark: 3
The autoregressive SaShiMi is originally evaluated using a form of rejection sampling [6] to retain only high probability generated samples for evaluation. We could also perform sampling tricks on all the models (since both diffusion and GAN models also have tractable likelihood measures). However, in the interest of an fair and simple comparison of the inherent performance of each model, we opt to keep the same and most general sampling method for each model (including ASGAN). So, in all experiments, we perform direct sampling from the latent space for the GAN and diffusion models according to the original papers. For the autoregressive models, we directly sample from the predicted distribution for each time-step.
V Results: Unconditional Speech Synthesis
V-A Comparison to baselines
TABLE I: Results measuring the quality and diversity of generated samples from unconditional speech synthesis models together with train/test set toplines for the SC09 dataset. Subjective MOS values with 95% confidence intervals are shown. IS, mIS, FID, AM are averaged over 5000 generated utterances generated from each model (or drawn from the train/test set for topline scores).
Model
IS\ $\uparrow$ mIS \ $\uparrow$ FID\ $\downarrow$ AM\ $\downarrow$ eMOS\ $\uparrow$ MOS\ $\uparrow$
Train set $9.37$ $237.6$ 00 $0.20$ $2.41$ $3.74\pm 0.12$
Test set $9.36$ $242.3$ $0.01$ $0.20$ $2.43$ $3.88\pm 0.12$
WaveGAN [1]
$4.45$ $34.6$ $1.77$ $0.81$ $1.06$ $2.88\pm 0.16$
DiffWave [5]
$5.13$ $49.6$ $1.68$ $0.68$ $1.66$ $3.43\pm 0.14$
SaShiMi [6]
$3.74$ $18.9$ $2.11$ $0.99$ $1.58$ $3.19\pm 0.15$
SaShiMi+DiffWave
$5.44$ $60.8$ $1.01$ $0.61$ $1.89$ $3.33\pm 0.12$
ASGAN (mel-spec.)
$7.02$ $162.8$ $0.56$ $0.36$ $1.76$ $3.51\pm 0.13$
ASGAN (HuBERT)
$7.67$ $226.7$ $0.14$ $0.26$ $1.99$ $3.68\pm 0.13$
TABLE II: Latent-space disentanglement and speed metrics. Speed is measured as the number of samples that can be generated per unit time on a single NVIDIA Quadro RTX 6000 using a batch size of 1, given in ksamples/sec. Some models do not have a $W$-space (WaveGAN) or any continuous latent space (SaShiMi).
Path length $\downarrow$ Separability $\downarrow$
Model
$Z$ $W$ $Z$ $W$ Speed $\uparrow$
WaveGAN [1]
1.03 4.86 2214.71
DiffWave [5]
[…]
SaShiMi [6]
0.14
SaShiMi+DiffWave
[…]
ASGAN (mel-spec.)
[…]
ASGAN (HuBERT)
[…]
We present our headline results in Table I, where we compare previous state-of-the-art unconditional speech synthesis approaches to the proposed ASGAN model. As a reminder, IS, mIS, FID and AM measure generated speech diversity and quality relative to the test set; eMOS and MOS are measures of generated speech naturalness. We see that both variants of ASGAN outperform the other models on most metrics. The HuBERT variant of ASGAN in particular performs best across all metrics. The improvement of the HuBERT ASGAN over the mel-spectrogram variant is likely because the high-level HuBERT speech representations make it easier for the model to disentangle common factors of speech variation. The previous best unconditional synthesis model, SaShiMi+DiffWave, still outperforms the other baseline models, and it appears to have comparable naturalness (similar eMOS and MOS) to the mel-spectrogram ASGAN variant. However, it appears to match the test set more poorly than either ASGAN variant on the other diversity metrics.
The latent space disentanglement metrics and generation speed are given in Table II. These results are more mixed, with WaveGAN being the fastest model and the one with the shortest latent path length in the $Z$-space. However, this is somewhat misleading since WaveGAN’s samples have low quality (naturalness, as measured by eMOS/MOS) and poor diversity compared to the other models (Table I). This means that WaveGAN’s latent space is a poor representation of the true distribution of speech in the SC09 dataset, allowing it to have a very small path length as most paths do not span a diverse set of speech variation.
In terms of linear separability, ASGAN again yields substantial improvements over existing models. The results confirm that ASGAN has indeed learned a disentangled latent space – a primary motivation for the model’s design. Specifically, this shows that the idea from image synthesis of using the latent $\mathbf{w}$ vector to linearly modulate convolution kernels can also be applied to speech. This level of disentanglement allows ASGAN to be applied to tasks unseen during training, evaluated later in Sec. VIII. Regardless of performance, the speed of all the convolutional GAN models (WaveGAN and ASGAN) is significantly better than the diffusion and autoregressive models, as reasoned in Sec. IV-B.
V-B Ablation experiments
TABLE III: Ablations of model design choices, comparing quality, diversity, and latent-space disentanglement of several ASGAN variants.
Quality & diversity Path length $\downarrow$
Variant
mIS \ $\uparrow$ FID\ $\downarrow$ eMOS\ $\uparrow$ $Z$ $W$
ASGAN (HuBERT)
$226.7$ $0.14$ $1.99$ $35.0$ $18.4$
w/o adaptive $D$ updates
$1.4$ $19.27$ $0.69$ $10.8$ $7.6$
w/o adaptive $D$ augmentation
$2.3$ $13.72$ $0.68$ $8.4$ $5.0$
w/o anti-aliasing filters
$102.7$ $3.31$ $1.81$ $75.8$ $63.1$
w/o modulated convolution
$2.4$ $10.96$ $0.63$ $111.4$ $48.5$
While the previous comparisons demonstrated the overall success of ASGAN’s design, we are still not certain which specific decisions from Sec. III are responsible for its performance. So, we perform several ablations of the HuBERT ASGAN model with specific components removed from the full model. Concretely, we ablate four key design choices: we train a variant without adaptive discriminator updates (Sec. III-D4), a variant without adaptive discriminator augmentation (Sec. III-D4), and a variant without any anti-aliasing filters (Sec. III-A3). Finally, we also train a variant without modulated convolutions (Sec. III-A2) such that $\mathbf{w}$ only controls the initial features passed to the generator’s convolutional encoder.
Table III shows the results for the ablated ASGAN approaches on a subset of the metrics. We see that on the quality and diversity metrics, the base ASGAN is best, while the models without adaptive discriminator updates and augmentation have better latent space disentanglement. However, recall that the main reason for including these was not to optimize disentanglement, but rather to ensure training stability and performance. As reasoned in Sec. III, without either adaptive updates or augmentation, the discriminator has a much easier task and begins to dominate the generator, confidently distinguishing between real and generated samples. So, while this makes optimization easier (leading to a smoother latent space), it means that the generator does not effectively learn from the adversarial task. A similar phenomenon can be seen with WaveGAN in Table II, where it scored well on the disentanglement metrics but had poor output quality in Table I.
When anti-aliasing filters are removed, both latent space disentanglement and synthesis quality are reduced, being slightly worse in all metrics compared to the full model. This validates our design motivation for the inclusion of low-pass filters to suppress aliased high-frequency content in the layer activations in Fig 1. Finally, the variant without the linear influence of $\mathbf{w}$ on each layer’s activations (i.e. without the modulated convolutions) is also worse than the baseline model in all metrics considered.
Overall, we can see from Table III that each of the key design aspects from Sec. III are necessary to achieve both high latent space disentanglement and synthesis quality in a single model – the main requirement for it to be performant on unseen downstream tasks, which we look at next.
VI Solving Unseen Tasks Through Linear Latent Operations
We have now shown intrinsically that ASGAN leads to a disentangled latent space. In this section and the ones that follow we show that ASGAN can also be used extrinsically to perform tasks unseen during training through linear manipulations of its latent space. From this point onwards we the HuBERT ASGAN variant.
As a reminder, the key aspect of ASGAN’s design is the latent space associated with the vector $\mathbf{w}$ is linearly disentangled. The idea is that, since the $\mathbf{w}$ vector can only linearly affect the output of the model through the Fourier feature layer and modulated convolutions (Fig. 1), $W$ must learn to linearly disentangle common factors of speech variation. If this turns out to be true – i.e. the space is indeed disentangled – then these factors should correspond to linear directions in the $W$ latent space. As originally motivated in studies on image synthesis [12, 13], this would mean that linear operations in the latent space should correspond to meaningful edits in the generated output. In our case, this would mean that if we know the relationship between two utterances, then the linear distance between the latent vectors $\mathbf{w}$ of those utterances should reflect that relationship. E.g. consider a collection of utterances differing only in noise level but otherwise having the same properties. If the space disentangles noise levels, then we should expect the latent points of these utterances to lie in the same linear subspace, reflective of the level of noise.
VI-A Projecting to the latent space
Before defining how unseen tasks can be phrased as latent operations, we first need to explain how we invert a provided utterance to a latent vector $\mathbf{w}$. We use a method similar to [13] where we optimize a $\mathbf{w}$ vector while keeping $G$ and the speech feature sequence $X$ fixed. Concretely, $\mathbf{w}$ is initialized to the mean $\bar{\mathbf{w}}=\mathbb{E}[W(\mathbf{z})]$ vector over 100k samples and then fed through the network to produce a candidate sequence $\tilde{X}$. An $L_{2}$ loss is then formed as the mean square distance between each feature in the candidate sequence $\tilde{X}$ and the target sequence $X$. Optimization follows [13], using Adam for 1000 iterations with a maximum learning rate of 0.1, and with Gaussian noise added to $\mathbf{w}$ in the first 750 iterations. The variance of this noise is set to be proportional to the average square $L_{2}$ distance between the mean $\bar{\mathbf{w}}$ and the sampled $\mathbf{w}$ vectors.
VI-B Downstream tasks
We look at several downstream tasks, each of which can be phrased as linear operations in the latent $W$-space.
VI-B1 Style mixing
We can perform voice conversion or speech editing by using style mixing [12] of the latent vectors $\mathbf{w}$. Concretely, we can project two utterances $X_{1}$ and $X_{2}$ to their latent representations $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$. Then – recalling the architecture in Fig. 1 – we can use different $\mathbf{w}$ vectors as input into each Style Block. According to our design motivation of the anti-aliasing filters in Sec. III-D5, the coarse styles are captured in the earlier layers, and the fine styles are introduced in later layers. So we can perform voice conversion from $X_{1}$’s speaker to $X_{2}$’s speaker by conditioning later layers with $\mathbf{w}_{2}$ while still using $\mathbf{w}_{1}$ in earlier layers. This causes the generated utterance to inherent the speaking style (fine) from the target utterance $X_{2}$, but retain the word identity (coarse) from $X_{1}$. By doing the opposite we can also do speech editing: having the speaker from $X_{1}$ say the word in $X_{2}$ by conditioning earlier layers with $\mathbf{w}_{2}$ while keeping $\mathbf{w}_{1}$ in later layers. Furthermore, because the $W$ latent space is continuous, we can interpolate between retaining and replacing the coarse and fine styles to achieve varying degrees of voice conversion or speech editing. We do these style mixing experiments in Sec. VIII-A.
VI-B2 Speech enhancement
Speech enhancement is the task of removing noise from an utterance [44]. Intuitively, if ASGAN’s $W$-space is linearly disentangled, there should be a single direction corresponding to increasing or decreasing the background noise in an utterance. Given several utterances only varying in degrees of noise, we can project them to the $W$-space and compute the direction in which to move to change the noise level. Concretely, to denoise an utterance $X_{0}$, we can generate $N$ additional utterances by adding increasingly more Gaussian noise, providing a list of utterances $X_{0},X_{1},...,X_{N}$, with $X_{n}=X_{0}+\mathcal{N}(\mathbf{0},n\sigma^{2}\mathbf{I})$. We then project each utterance to the latent space, yielding $\mathbf{w}_{0},\mathbf{w}_{1},...,\mathbf{w}_{N}$. To get a single vector corresponding to the direction of decreasing noise, we compute the average unit vector from the higher-noise vectors to the original latent vector:
$\boldsymbol{\delta}=\frac{1}{N}\sum_{n=1}^{N}\frac{\mathbf{w}_{0}-\mathbf{w}_{n}}{||\mathbf{w}_{0}-\mathbf{w}_{n}||_{2}}$
Now we can denoise the original utterance $X_{0}$ by moving in the direction of $\boldsymbol{\delta}$ in the latent space. In Sec. VIII we evaluate this method and investigate how far we can move in the $\boldsymbol{\delta}$-direction.
VI-B3 Speaker verification and keyword classification
The previous downstream tasks were generative in nature. The $W$ latent space also allows us to perform discriminative tasks such as speaker verification and keyword classification: determining which speaker or word is present in a given utterance, respectively. For both tasks we use the linear nature of disentanglement: given enrollment utterances containing labelled speakers (for speaker verification) or words (for keyword classification), we invert them to their $\mathbf{w}$ vectors. Then we find the linear projection within the $W$ latent space that maximizes the separation of the labeled characteristic using linear discriminant analysis (LDA). For inference on new data, we invert the input, project it along the LDA axes, and make a decision based on linear distances to other points in the LDA-projected latent space.
In speaker verification, we need to predict a score corresponding to whether two utterances are spoken by the same speaker [45]. The speakers are both unseen during training. So, we project all enrollment and test utterances to the $W$ latent space and then to the LDA axes (with axes fit on training data to maximize speaker separation). The speaker similarity score is then computed as the cosine distance between two vectors in this space. For evaluation, these scores can be used to compute an equal error rate (EER) [45]. Similarly, for keyword classification, we project everything to the LDA axes maximizing content (i.e. what digit is spoken), compute the centroids of points associated with each digit (since we know the word labels beforehand), and then assign new utterances and their corresponding projected points to the label of the closest centroid by cosine distance. We compare our latent-space LDA-based speaker verification and keyword classification approaches to task-specific models in Sec. VIII.
VII Experimental Setup: Unseen Tasks
VII-A Evaluation metrics
To evaluate ASGAN’s performance in each unseen task, we use the standard objective metrics from the literature:
  • Voice conversion: We measure conversion intelligibility following [46, 47], whereby we perform voice conversion and then apply a speech recognition system to the output and compute a character error rate (CER) and $F_{1}$ classification score to the word spoken in the original utterance. Speaker similarity is measured as described in [46] whereby we find similarity scores between real/generated utterance pairs using a trained speaker classifier, and then compute an EER with real/generated scores assigned a label of 0 and real/real pair scores assigned a label of 1.
  • Speech enhancement: Given a series of original clean and noisy utterances, and the models’ denoised output, we compute standard measures of denoising performance: narrow-band perceptual evaluation of speech quality (PESQ) [48] and short term objective intelligibility (STOI) scores [49].
  • Speaker verification: Using the similarity scores on randomly sampled pairs of utterances from matching and non-matching speakers, we compute an EER as the measure of performance. We pair each evaluation utterance with another utterance from the same or different speaker with equal probability.
  • Keyword classification: We use the standard classification metrics of accuracy and $F_{1}$ score.
Figure 2: Voice conversion interpolation in the $W-$latent space. Given a source (top left) and target utterance (top right), we smoothly convert the speaker from the source to the reference by linearly interpolating the projected $\mathbf{w}$ vectors from $\mathbf{w}_{1}$ (source) to $\mathbf{w}_{2}$ (target) for use in the fine styles. The $W$-space interpolation is illustrated as a 2D linear discriminant analysis (LDA) decomposition of the SC09 test set.
VII-B Baseline systems
For each downstream task, we compare ASGAN to a strong task-specific baseline trained on the SC09 dataset. For voice conversion, we compare against AutoVC [50] – a well-known voice conversion model. Specifically, we use the model and training setup defined in [50] and trained on the SC09 dataset [^4] https://github.com/RF5/simple-autovc. For speech enhancement, we compare to MANNER [51], a recent high-performing speech enhancement model operating in the time-domain. Since we have no paired clean/noisy utterances for the SC09 dataset, we follow the technique from [52] to construct a speech enhancement dataset. Concretely, we assume each utterance in the SC09 dataset is a clean utterance and randomly add additional Gaussian noise to each utterance corresponding to a signal-to-noise ratio of 0 to 10 dB, sampled uniformly. We then train MANNER using the settings and code provided by the original authors [^5] https://github.com/winddori2002/MANNER on this constructed dataset. For speaker verification, we use an x-vector model from [53], trained using the code and optimization settings from an open-source implementation [^6] https://github.com/RF5/simple-speaker-embedding on the SC09 training set, taking the final checkpoint with best validation performance. Finally, for keyword classification we, use the same ResNeXT classifier described in Sec. IV-B.
VIII Results: Unseen Tasks
We apply ASGAN to a range of tasks that it didn’t see during training, comparing it to established baselines. Note, however, that the goal isn’t to outperform these tailored systems on every task, but rather to show that a single model, ASGAN, can provide robust performance across a range of tasks that it was never trained on.
VIII-A Voice conversion
Following Sec. VI-B1, we use ASGAN for voice conversion. We compare its performance to that of AutoVC (Sec. VII-B). To test the models, we sample reference utterance from a different target speaker for each utterance in the SC09 test set, yielding 4107 utterance pairs on which we perform inference. Through initial validation experiments, we found that the optimal interpolation amount for the fine styles was to set the input to the later Style Blocks as $\mathbf{w}_{1}+1.75(\mathbf{w}_{2}-\mathbf{w}_{1})$ for $\mathbf{w}_{1}$ from the source utterance and $\mathbf{w}_{2}$ from the reference.
The results are shown in Table IV. In terms of similarity to the target speaker, AutoVC is superior to ASGAN, while ASGAN is superior in terms of intelligibility. Recall that our goal is not to outperform specialized models trained for specific tasks, but to demonstrate that ASGAN can generalize to diverse tasks unseen during training. While not superior to AutoVC, the scores in Table IV are competitive in voice conversion.
To give better intuition to the linear nature of the latent space, we show an example of interpolating the fine styles smoothly from one speaker to another in Fig. 2: we project the latent $\mathbf{w}$ points from many utterances in the test set using LDA fit on the speaker label. The figure raises two interesting observations. First, we see that speaker identity is largely a linear subspace in the latent space because of the strong LDA clustering observed. Second, as we interpolate from one point to another in the latent space, the intermediary points are still intelligible and depict realistic mixing of the source and target speakers, while leaving the content unchanged. We encourage the reader to listen to audio samples at https://rf5.github.io/slt2022-asgan-demo.
TABLE IV: Unseen generative task performance of ASGAN compared to task-specific systems (AutoVC [50] for voice conversion, MANNER [51] for speech enhancement).
Voice conversion (%) Speech enhancement
Model
EER \ $\uparrow$ CER \ $\downarrow$ F1 \ $\uparrow$ PESQ\ $\uparrow$ STOI\ $\uparrow$
Task-specific baseline $32.3$ $64.8$ $29.1$ $2.13$ $81.4$
ASGAN (HuBERT)
$29.7$ $37.6$ $61.3$ $0.95$ $21.2$
VIII-B Speech enhancement
[…]
Following Sec. VI-B1, we use ASGAN for speech enhancement, with $N=10$, $\sigma^{2}=0.001$. Using the process described in Sec. VII-B, we evaluate both ASGAN and the task-specific MANNER system on the SC09 test set in Table IV. From the results we see that ASGAN performs worse than MANNER. As was the case in voice conversion, this isn’t unexpected, given that MANNER is a specialized task-specific model, trained specifically for speech enhancement. Meanwhile, ASGAN is only trained to accurately model the density of utterances in the SC09 dataset. The performance of ASGAN on this task indicates some level of denoising. We suspect some of the performance limitations are due to the relatively simple inversion method of Sec. VI-A, which allows for only a crude approximation of the ideal vector $\mathbf{w}$ corresponding to a speech utterance $X$. Using a more sophisticated latent projection method such as pivot tuning inversion [54] would likely reduce the discrepancy between $\tilde{X}$ and $X$. Our goal here, however, was simply to illustrate ASGAN’s denoising capability without any additional fine-tuning of ASGAN’s generator.
[…]
VIII-C Speaker verification
TABLE V: Unseen discriminative task performance of ASGAN compared to task-specific systems (GE2E RNN [53] for speaker verification, ResNeXT classifier [38] for keyword classification).
Speaker verification Keyword classification
Model
EER \ $\downarrow$ Accuracy \ $\uparrow$ F1 \ $\uparrow$
Task-specific baseline $8.43$ $98.20$ $98.20$
ASGAN (HuBERT)
$41.54$ $72.58$ $72.78$
ASGAN (HuBERT) + LDA
$30.96$ $90.82$ $90.82$
Figure 4: Keyword classification and speech editing in the $W$-latent space. Given a source (top left) and target utterance (top right), we smoothly convert the content from the source to the reference by linearly interpolating the projected $\mathbf{w}$ vectors from $\mathbf{w}_{1}$ (source) to $\mathbf{w}_{2}$ (target) for use in the coarse styles. The $W$-space interpolation is illustrated as a 2D LDA decomposition on the SC09 test set fit on spoken digits, showing its usefulness for keyword classification.
We next consider a discriminative task: speaker verification, as outlined in Sec. VI-B3. Using the evaluation protocol of Sec. IV-B, we obtain the results in Table V on the SC09 test set. The first ASGAN entry performs the naive cosine comparison, while the ‘+ LDA’ entry compares distances after first applying LDA as described in Sec. VI-B3. As with the generative tasks above, we see that ASGAN is competitive, but not superior to the task-specific GE2E model [53]. The naive comparison in the regular latent space achieves fairly poor results. However, comparing the naive comparison method to the LDA projection approach, we see an over 10% EER improvement, with ASGAN achieving an EER of 30.96% despite never seeing any of the speakers in the test set. This reinforces the observations from Fig. 2, where we can see that the LDA projection of latent $\mathbf{w}$ vectors disentangles speaker identity from other aspects of speech. And, because LDA is a linear operation, we have not discarded any content information. The speaker verification performance could likely be improved if a deep model was trained to cluster speakers in the latent space. But, again, our goal is not state-of-the-art performance on this task, but to show that ASGAN can achieve reasonable accuracy on a task for which it was not trained—owing to its linearly disentangled latent space.
VIII-D Keyword classification
Finally we apply ASGAN to keyword classification. We fit LDA on digit spoken using the SC09 validation set as enrollment, and compute predictions following Sec. VI-B3. The results compared to the ResNeXT classifier are given in Table V using both the naive and LDA comparison method. As before, we see a large performance jump when using the LDA projection, with the best ASGAN method only performing roughly 7% worse than the specialized classification model. A graphical illustration of how word classes are disentangled through LDA is given in Fig. 4. We show both the LDA projection of latent points (colored by their ground truth labels) and also how this disentanglement can be used to perform speech editing using the fine styles described in Sec. VI-B1, where we replace the word spoken but leave the speaker identity and other characteristics of the utterance intact. It allows us to smoothly interpolate between one spoken word and another. This could potentially be useful, e.g., for perceptual speech experiments. From both the LDA plot in Fig. 4 and Table V, we again observe strong evidence that our design for disentanglement is successful.
IX Conclusion
We introduced ASGAN, a model for unconditional speech synthesis designed to learn a disentangled latent space. We adapted existing and incorporated new GAN design and training techniques to enable ASGAN to learn a continuous, linearly disentangled latent space in order to outperform existing autoregressive and diffusion models. Experiments on the SC09 dataset demonstrated that ASGAN outperforms previous state-of-the-art models on most unconditional speech synthesis metrics, while also being substantially faster. Further experiments also demonstrated the benefit of the disentangled latent space: ASGAN can, without any additional training, perform several speech processing tasks in a zero-shot fashion through linear operations in its latent space, showing reasonable performance in voice conversion, speech enhancement, speaker verification, and keyword classification.
One major limitation of our work is scale: once trained, ASGAN can only generate utterances of a fixed length, and the model struggles to generate coherent full sentences on datasets with longer utterances (a limitation shared by existing unconditional synthesis models [1, 5, 6]). Furthermore, the method used to project utterances to the latent space could be improved by incorporating recent inversion methods [54]. Future work will aim to address these shortcomings.